Problem 6.22 6 of 21 (> Con PartA How many Ar+ ions are in a mole of AlBry? Express your answer using four significant figures. Submit Request Answer Part B How many Br ions are in a mole of AlBr3? Express your answer using four significant figures. Submit Request Answer Provide Feedback O Type here to search
Answer #1

Part A:

1 mole AlBr3 has one mole of Al3+ ions and 3 moles of Br- ions.

N = 6.022 * 1023 ions of Al3+

Part B:

1 mole AlBr3 has one mole of Al3+ ions and 3 moles of Br- ions

N = 3 * 6.022 * 1023 ions of Br- = 1.807 * 1024 ions of Br-